# Low-Overhead Channel Estimation via 3D Extrapolation for TDD mmWave Massive MIMO Systems Under High-Mobility Scenarios

### v1

@@ II-B Problem Formulation
To support the measurement of all transmit ports and subcarriers, the length of the pilot signal is $N_{R}\times N_{c}$, Therefore, the pilot training overhead is $(N_{RF}\times\frac{N_{T}}{N_{RF}})\times N_{R}\times N_{c}=N_{T}\times N_{R}\times N_{c}$. Considering a large number of antennas and subcarriers in the system, the pilot training overhead is extremely high and significantly challenging to the system’s spectral efficiency. In addition, the time and power consumption due to the switching of antennas connected to RF chains is also unaffordable. Moreover, the pilot signal should be transmitted for uplink channel estimation every channel coherence time $T_{c}$ to avoid significant channel aging. For simplicity, we call this scheme the general pilot training scheme hereafter. We denote the pilot signal period as $T_{p}$ and for the general pilot training scheme we have $T_{p}\leq T_{c}$. Note that as specified in the 3GPP 5G technical specification [21], the frame structure in the 5G new radio (NR) is shown in Fig. 1 (c). Each frame spans $10$ ms and is divided into two half-frames. Each half-frame spans $5$ ms and contains five sub-frames, with each sub-frame spanning $1$ ms. The number of slots in one sub-frame, i.e., $N_{slot}$, is determined by the numerology $\mu$, and $N_{slot}=2^{\mu}$. Here we assume $\mu=3$ which indicates $N_{slot}=8$ in one sub-frame and each slot spans $0.125$ ms. It should be emphasized that under scenarios with high carrier frequency and high UE mobility, the channel coherence time could be very short. For instance, when operating at a carrier frequency of $28$ GHz and with the UE velocity of $60$ km/h, the channel coherence time is approximately $0.32$ ms[18]. With the general pilot training scheme, the TDD slot pattern can be illustrated by Fig. 1 (d), where ‘S’ stands for a special slot configured for uplink channel estimation and ‘D’ stands for a downlink slot for downlink data transmission. [^2] Note that for ease of elaboration, we do not consider uplink data transmission slots in the sub-frame. Our method can be easily extended to sub-frames with uplink data transmission slots. The pilot signal has to be transmitted every two slots (i.e., $T_{p}=0.25$ ms) to mitigate channel aging with $T_{c}=0.32$ ms, further degrading the system’s spectral efficiency. [^3] Note that in Fig. 1 (d), (e), and (f), we assume the pilot symbol is configured at the last symbol index of each special slot.
Defining the single-slot pilot training overhead $C_{sl}$ as the pilot training overhead for one-time uplink channel estimation (i.e., the spatial-frequency domain pilot overhead), the overall pilot training overhead $C_{o}$ for an interval of length $T$ is defined as
@@
From the previous discussions, the overall pilot training overhead $C_{o}$ is huge, since the single-slot pilot training overhead $C_{sl}$ is large due to the high-dimensional CSI matrices and the pilot signal period $T_{p}$ is short in high-mobility scenarios. To circumvent such huge overall pilot training overhead, we consider spatial, frequency, and temporal channel extrapolation to systematically reduce the pilot training overhead from two perspectives: 1) reducing the single-slot pilot training overhead $C_{sl}$ (i.e., the spatial-frequency domain pilot overhead) via spatial-frequency channel extrapolation; 2) enlarging the pilot signal period $T_{p}$ via slot-level channel extrapolation and thereby reducing the temporal domain pilot overhead. In addition, it is worth noting that through spatial extrapolation, the substantial time and power consumption due to antenna switching can be avoided. And through the slot-level channel extrapolation, more slots can be configured for data transmission to improve the system’s spectral efficiency. The overall framework of the proposed method is shown in Fig. 2.
@@
On the other hand, we propose to enlarge the pilot signal period $T_{p}$ by leveraging the uplink-downlink channel reciprocity and temporal dependencies for slot-level channel extrapolation (SLCE). In TDD systems, it is commonly assumed that the uplink and downlink channels are reciprocal during the channel coherence time, such that the downlink channel can be derived from the estimated uplink channel. Note that due to high UE mobility and short channel coherence time, the uplink-downlink channel reciprocity only holds for the special slot of a sub-frame and a few downlink data transmission slots in the same sub-frame. And due to the hardware asymmetry in transceivers, uplink-downlink channel calibration is required to derive the downlink channel accurately from the estimated uplink channel. Moreover, like many real-world phenomena, wireless channels exhibit temporal dependencies over time (e.g., explicit patterns like trends, etc). Capturing these temporal dependencies from downlink channels allows us to conduct accurate temporal extrapolation for future slots. However, the temporal dependencies for high-dimensional channels are challenging to capture. Fortunately, with DNNs, it is promising to learn the uplink-downlink channel reciprocity and temporal dependencies for SLCE. Then, with SLCE, the pilot signal can be transmitted with a pilot signal period $T_{p}\gg T_{c}$. Taking transmitting the pilot signal only once per sub-frame and at the first slot of each sub-frame as an example, the TDD slot pattern with the SLCE-aided pilot training scheme is provided in Fig. 1 (e). [^4] Note that for ease of elaboration, in the following context, we simply use this TDD slot pattern and suppose the pilot signal only once per sub-frame. The pilot signal can be set flexibly in practice. In this example, the pilot signal period is $T_{p}=1$ ms, which is much larger than the channel coherence time $T_{c}=0.32$ ms. To tackle the channel aging issue caused by high carrier frequency and high UE mobility, SLCE is thus anticipated to predict the downlink channels within the pilot signal period accurately for downlink data transmission. Specifically, based on the estimated uplink channel at the special slot, the downlink channel at the first downlink data transmission slot can be initially acquired by exploiting the uplink-downlink channel reciprocity and via channel calibration. Then, by learning the temporal dependencies among time-varying downlink channels, accurate estimates of the downlink channels at all other downlink data transmission slots can be achieved. Therefore, denoting $\mathbf{H}^{d}\{s,t\}\in\mathbb{C}^{N_{R}\times(N_{T}\times N_{c})}$ as the downlink channel at the $t$-th slot of the $s$-th sub-frame, the SLCE problem can be formulated as
@@ II-C Sounding Reference Signal Pattern
In the following two sections, to reduce the overall pilot training overhead $C_{o}$ from the aforementioned two perspectives, i.e., reducing the single-slot pilot training overhead $C_{sl}$ and enlarging the pilot signal period $T_{p}$, we propose the KDD-SFCEN for single-slot uplink channel estimation and the TUDCEN for slot-level channel extrapolation respectively.
@@ III Spatial-Frequency Channel Extrapolation based Single-Slot Uplink Channel Estimation
Considering the marginal effects of pilots in the spatial and frequency domains, we propose to conduct joint spatial-frequency channel extrapolation to reduce the pilot training overhead of single-slot uplink channel estimation. Traditional channel estimation methods based on domain knowledge have been fully validated for decades that they can provide feasible channel estimates in time-invariant channels, which can be exploited as initial estimates for deep learning-based channel estimation methods. In addition, to estimate the uplink channel accurately with only a few pilots in both the spatial and frequency domains, the spatial and frequency characteristics of massive MIMO channels must be fully exploited [11]. Therefore, in this section, we propose the KDD-SFCEN for single-slot uplink channel estimation in TDD massive MIMO systems, which benefits from both domain knowledge via the traditional knowledge-driven channel estimation and data via joint spatial-frequency channel extrapolation. As shown in Fig. 2, the proposed network consists of two components, i.e., a knowledge-driven coarse channel estimator and a spatial-frequency channel extrapolator, for coarse channel estimation and spatial-frequency channel extrapolation, respectively.
@@ III-A Coarse Channel Estimation
Since traditional LS estimation has been demonstrated to be a simple but effective method to achieve coarse channel estimates, we exploit this domain knowledge and use the traditional LS estimator as the knowledge-driven coarse channel estimator to accelerate the training of our proposed uplink channel estimation network. Specifically, the received pilot signal $\mathbf{Y}^{p}{\{s,0\}}$ and the corresponding transmitted diagonal pilot signal $\mathbf{S}^{p}{\{s,0\}}$ is first processed by an LS estimator as
@@ III-B Spatial and Frequency Channel Extrapolation
To estimate the uplink channel accurately with a few pilots, we propose the attention-based sub-element extrapolation module (ASEEM) to fully exploit the spatial and frequency characteristics of massive MIMO channels. In addition, directly extrapolating partial CSI to full CSI is significantly challenging, especially when the spatial and frequency compression ratios are large. Therefore, we propose to conduct spatial and frequency channel extrapolation progressively. As Fig. 2 shows, the proposed spatial-frequency channel extrapolator consists of a progressive spatial extrapolation block and a progressive frequency extrapolation block for spatial and frequency extrapolation, respectively. The progressive spatial extrapolation block and the progressive frequency extrapolation block consist of $N_{SE}$ spatial ASEEMs and $N_{FE}$ frequency ASEEMs, respectively. These spatial/frequency ASEEMs progressively extrapolate partial spatial/frequency CSI to full spatial/frequency CSI.
@@ III-C Model Training
After obtaining the estimated uplink channel, the slot-level channel extrapolation can then be conducted to enlarge the pilot signal period $T_{p}$ and further reduce the overall pilot training overhead $C_{o}$.
@@ IV Channel Reciprocity and Temporal Dependency Aided Slot-Level Channel Extrapolation
An accurate initial state is essential to the subsequent extrapolations in the temporal domain. Thanks to the uplink-downlink channel reciprocity in TDD systems, the downlink channel of the first downlink slot of a sub-frame can be accurately derived from the uplink channel of the sub-frame’s special slot via uplink-downlink channel calibration. The derived downlink channel provides an accurate initial state for predicting downlink channels at future downlink slots. Recent advances in generative AI, e.g., the ChatGPT[28], have witnessed the superiority of autoregressive generative models, specifically generative Transformers[20], for processing sequential data. For time-varying channels, which are also one kind of sequential data, generative Transformers are also promising on account of their capabilities in learning complicated temporal dependencies and autoregressive modeling. Therefore, in this section, we propose the TUDCEN consisting of the UDCCN and the DCEN to achieve slot-level channel extrapolation. The architecture of the TUDCEN is shown in Fig. 2. Specifically, the UDCCN realizes uplink-downlink channel calibration for initialization by exploiting the uplink-downlink channel reciprocity, and the DCEN mainly exploits the proposed spatial-frequency sampling embedding module to reduce the computational complexity and generative Transformers for temporal dependency learning and autoregressive downlink channel extrapolation.
@@ IV-A Channel Reciprocity Aided Uplink-Downlink Channel Calibration
Although the physical uplink-downlink channels within the channel coherence time are generally reciprocal, uplink-downlink channel calibration is practically necessary to compensate for the hardware asymmetry in transceivers[22]. The uplink-downlink channel extrapolation can be formulated by
@@ IV-B Temporal Dependency Learning and Downlink Channel Extrapolation
It is worth mentioning that the downlink channel extrapolation problem is conducted along the time domain and aims to predict channels at future slots, which can be regarded as a time series prediction problem. However, different from general time series prediction problems, historical observations in the investigated downlink channel extrapolation problem (i.e., downlink channels at all downlink slots) can not be obtained unless huge time-frequency resources were spent to conduct channel estimation for all downlink slots. As a result, the performance of widely adopted sequence-to-sequence (Seq2Seq) prediction schemes in related works[18] would be significantly degraded due to the lack of available historical channel information. Note that like many real-world phenomena, wireless channels exhibit temporal dependencies over time. Then, based on the chain rule in the probability theory, the joint probability of the concatenated downlink channels can be expressed as
@@
where $F_{dce}$ denotes the DCEN, such that (IV-B) can be achieved and the downlink channels from the $2$nd slot to the ($N_{slot}-1$)-th slot can be extrapolated given only $\hat{\mathbf{H}}^{d}\{s,1\}$. It is worth noting that (IV-B) actually refers to the autoregressive generation manner and the DCEN should thus be an autoregressive model. And considering that $\hat{\mathbf{H}}^{d}\{s,t\},t=1,2,\ldots,N_{slot}-1$ are dependent, capturing these temporal dependencies among them would be beneficial to achieve (IV-B). To this end, we propose the DCEN constituted by (inverse) spatial-frequency patch embedding layers and generative Transformers for temporal dependency learning and autoregressive downlink channel extrapolation. The architecture of the DCEN is also shown in Fig. 2.

### v2

@@ II-B Problem Formulation
To support the measurement of all transmit ports and subcarriers, the length of the pilot signal is $N_{R}\times N_{c}$. Therefore, the pilot training overhead is $(N_{RF}\times\frac{N_{T}}{N_{RF}})\times N_{R}\times N_{c}=N_{T}\times N_{R}\times N_{c}$, which is extremely high and significantly challenging to the system’s spectral efficiency. In addition, the time and power consumption due to the switching of antennas connected to RF chains is also unaffordable. Moreover, the pilot signal should be transmitted for uplink channel estimation every channel coherence time $T_{c}$ to avoid significant channel aging. For simplicity, we call this scheme the general pilot training scheme hereafter. We denote the pilot signal period as $T_{p}$ and for the general pilot training scheme we have $T_{p}\leq T_{c}$. Note that as specified in the 3GPP 5G technical specification [22], the frame structure in the 5G new radio (NR) is shown in Fig. 1 (c). Each frame spans $10$ ms and is divided into two half-frames. Each half-frame spans $5$ ms and contains five sub-frames, with each sub-frame spanning $1$ ms. The number of slots in one sub-frame, i.e., $N_{slot}$, is determined by the numerology $\mu$, and $N_{slot}=2^{\mu}$. Here we assume $\mu=3$ which indicates $N_{slot}=8$ in one sub-frame and each slot spans $0.125$ ms. It should be emphasized that under scenarios with high carrier frequency and high UE mobility, the channel coherence time could be very short. For instance, when operating at a carrier frequency of $28$ GHz and with the UE velocity of $60$ km/h, the channel coherence time is approximately $0.32$ ms[18]. With the general pilot training scheme, the pilot signal has to be transmitted every two slots (i.e., $T_{p}=0.25$ ms) to mitigate channel aging with $T_{c}=0.32$ ms as Fig. 1 (d) shows, [^3] Note that for ease of elaboration, we do not consider uplink data transmission slots in the sub-frame. Our method can be easily extended to sub-frames with uplink data transmission slots. Additionally, in Fig. 1 (d), (e), and (f), we assume the pilot symbol is configured at the last symbol index of each special slot. further degrading the system’s spectral efficiency.
Denote the single-slot pilot training overhead $C_{sl}$ as the pilot training overhead for single-slot uplink channel estimation (i.e., the spatial-frequency domain pilot overhead), which can be mathematically expressed as
$C_{sl}=N_{T}\times N_{R}\times N_{c}.$ (7)
Then, for an interval of length $T$, the uplink channel estimation needs to be conducted for $\frac{T}{T_{p}}$ times, leading to huge temporal domain pilot overhead. Therefore, the overall pilot training overhead $C_{o}$ can be defined as
@@
To circumvent such huge overall pilot training overhead, we consider spatial, frequency, and temporal channel extrapolation to systematically reduce the pilot training overhead from two perspectives: 1) reducing the single-slot pilot training overhead $C_{sl}$ via spatial-frequency channel extrapolation; 2) reducing the times of single-slot channel estimations $\frac{T}{T_{p}}$ to reduce the temporal domain pilot overhead via temporal channel extrapolation. In addition, it is worth noting that through spatial extrapolation, the substantial time and power consumption due to antenna switching can be avoided. And through slot-level channel extrapolation, more slots can be configured for data transmission to improve the system’s spectral efficiency. The overall framework of the proposed method is shown in Fig. 2.
@@
On the other hand, we propose to reduce the times of single-slot channel estimations $\frac{T}{T_{p}}$ by leveraging the uplink-downlink channel reciprocity and temporal dependencies for slot-level channel extrapolation (SLCE). In TDD systems, it is commonly assumed that the uplink and downlink channels are reciprocal during the channel coherence time, such that the downlink channel can be derived from the estimated uplink channel. Note that due to high UE mobility and short channel coherence time, the uplink-downlink channel reciprocity only holds for the special slot of a sub-frame and a few downlink data transmission slots in the same sub-frame. And due to the hardware asymmetry in transceivers, uplink-downlink channel calibration is required to derive the downlink channel accurately from the estimated uplink channel. Moreover, like many real-world phenomena, wireless channels exhibit temporal dependencies over time (e.g., explicit patterns like trends, etc). Capturing these temporal dependencies from downlink channels allows us to conduct accurate temporal extrapolation for future slots. However, the temporal dependencies for high-dimensional channels are challenging to capture. Fortunately, with DNNs, it is promising to learn the uplink-downlink channel reciprocity and temporal dependencies for SLCE. Then, with SLCE, the pilot signal can be transmitted with a pilot signal period $T_{p}\gg T_{c}$. Taking transmitting the pilot signal only once per sub-frame and at the first slot of each sub-frame as an example, the TDD slot pattern with the SLCE-aided pilot training scheme is provided in Fig. 1 (e). [^4] Note that for ease of elaboration, in the following context, we simply use this TDD slot pattern and suppose the pilot signal only once per sub-frame. The pilot signal can be set flexibly in practice. In this example, the pilot signal period is $T_{p}=1$ ms, which is much larger than the channel coherence time $T_{c}=0.32$ ms. To tackle the channel aging issue caused by high carrier frequency and high UE mobility, SLCE is thus anticipated to predict the downlink channels within the pilot signal period accurately for downlink data transmission. Specifically, based on the estimated uplink channel at the special slot, the downlink channel at the first downlink data transmission slot can be initially acquired by exploiting the uplink-downlink channel reciprocity and via channel calibration. Then, by learning the temporal dependencies among time-varying downlink channels, accurate estimates of the downlink channels at all other downlink data transmission slots can be achieved. Therefore, denoting $\mathbf{H}^{d}\{s,t\}\in\mathbb{C}^{N_{R}\times(N_{T}\times N_{c})}$ as the downlink channel at the $t$-th slot of the $s$-th sub-frame, the SLCE problem can be formulated as
@@ II-C Sounding Reference Signal Pattern
In the following two sections, to reduce the overall pilot training overhead $C_{o}$ from the aforementioned two perspectives, i.e., reducing the single-slot pilot training overhead $C_{sl}$ and reducing the times of single-slot channel estimations $\frac{T}{T_{p}}$, we propose the KDD-SFCEN for single-slot uplink channel estimation and the TUDCEN for slot-level channel extrapolation respectively.
@@ III Spatial-Frequency Channel Extrapolation based Single-Slot Uplink Channel Estimation
Considering the marginal effects of pilots in the spatial and frequency domains, we propose to conduct joint spatial-frequency channel extrapolation to reduce the pilot training overhead of single-slot uplink channel estimation. Traditional channel estimation methods based on domain knowledge have been fully validated for decades that they can provide feasible channel estimates in time-invariant channels, which can be exploited as initial estimates for deep learning-based channel estimation methods. In addition, to estimate the uplink channel accurately with only a few pilots in both the spatial and frequency domains, the spatial and frequency characteristics of massive MIMO channels must be fully exploited [11]. Therefore, in this section, we propose the KDD-SFCEN for single-slot uplink channel estimation in TDD massive MIMO systems, which benefits from both domain knowledge via the traditional knowledge-driven channel estimation and data via joint spatial-frequency channel extrapolation. As shown in Fig. 2, the proposed network consists of a knowledge-driven coarse channel estimator and a spatial-frequency channel extrapolator for coarse channel estimation and spatial-frequency channel extrapolation, respectively.
@@ III-A Coarse Channel Estimation
Since traditional LS estimation has been demonstrated to be a simple but effective method to achieve coarse channel estimates, we exploit the traditional LS estimator as the knowledge-driven coarse channel estimator to accelerate the training of our proposed uplink channel estimation network. Specifically, the received pilot signal $\mathbf{Y}^{p}{\{s,0\}}$ and the corresponding transmitted diagonal pilot signal $\mathbf{S}^{p}{\{s,0\}}$ is first processed by an LS estimator as
@@ III-B Spatial and Frequency Channel Extrapolation
To estimate the uplink channel accurately with a few pilots, we propose the spatial-frequency channel extrapolator consisting of ASEEMs and the progressive extrapolation architecture to exploit spatial and frequency correlations of wireless channels progressively for spatial and frequency domain extrapolation. The spatial correlation refers to the correlation between the signals received at different antennas in the antenna domain, which practically depends on the spatial characteristics of the environment (e.g., multipath components, angle of arrival, etc.) and antenna configurations (e.g., antenna geometry, antenna spacing, etc). The frequency correlation refers to the correlation of the channel across different subcarriers in the frequency domain, which is mainly influenced by the channel’s multipath delay spread and Doppler shift due to mobility. In addition, it can also be observed that the closer the antennas/subcarriers are, the stronger the correlations among them are. Based on these observations, we design the progressive extrapolation architecture to first focus on capturing the correlations between neighboring antennas/subcarriers, and then progressively broaden the scope to capture correlations, with the ASEEM designed to capture specific spatial/frequency correlations in the process. As Fig. 2 shows, the proposed spatial-frequency channel extrapolator comprises a progressive spatial extrapolation block with $N_{SE}$ spatial ASEEMs and a progressive frequency extrapolation block with $N_{FE}$ frequency ASEEMs for spatial and frequency extrapolation, respectively. These spatial/frequency ASEEMs progressively extrapolate partial spatial/frequency CSI to full spatial/frequency CSI.
@@ III-C Model Training
Remark: It can be noticed that many works on mmWave channel estimation estimate sparse parameters of mmWave channels exploiting channel sparsity. However, we don’t intentionally introduce channel sparsity to design our 3D channel extrapolation framework. The reasons are two-fold. On the one hand, exploiting the sparsity would require some prior knowledge of the channels’ specific sparse representation (e.g., the angular spread of angle of arrival (AoA) and angle of departure (AoD)). Although mmWave channels generally exhibit sparsity, in some indoor mmWave communication scenarios with rich multipath components, it is difficult to obtain an accurate sparse representation of the channels, leading to degraded channel estimation performance of channel sparsity-based channel estimation methods. On the other hand, even in non-mmWave systems without channel sparsity, our 3D channel extrapolation framework is expected to work well to reduce the power consumption of RF chains. Therefore, channel sparsity is not incorporated into the design of the 3D channel extrapolation framework to generalize its application scenarios. Nonetheless, it is worth noting that for scenarios with explicit channel sparsity, the 3D channel extrapolation framework can be easily extended to exploit channel sparsity to further reduce the pilot overhead and computational complexity. Specifically, the 3D channel extrapolation framework is currently trained to estimate the full channel matrix given the supervision of the full channel matrix. Suppose prior knowledge of the channel’s sparse representation is available, the 3D channel extrapolation framework can be trained to first estimate sparse parameters and then reconstruct the full channel matrix with two-level supervision given labels of both sparse parameters and the full channel matrix.
After obtaining the estimated uplink channel, the slot-level channel extrapolation can then be conducted to reduce the times of uplink channel estimations $\frac{T}{T_{p}}$.
@@ IV Channel Reciprocity and Temporal Dependency Aided Slot-Level Channel Extrapolation
An accurate initial state is essential to temporal channel extrapolation. In TDD systems, the downlink channel of the first downlink slot of a sub-frame can be accurately derived from the uplink channel of the sub-frame’s special slot via uplink-downlink channel calibration. This provides an accurate initial state for predicting future downlink channels. Recent advances in generative artificial intelligence (GAI), such as ChatGPT[27], have demonstrated the power of autoregressive models like generative Transformers[21] in handling sequential data. Given that time-varying channels also form sequential data, generative Transformers are promising on account of their capabilities in learning complicated temporal dependencies and autoregressive modeling. Accordingly, we propose the TUDCEN, comprising the UDCCN and the DCEN, for slot-level channel extrapolation. The architecture of the TUDCEN is shown in Fig. 2. Specifically, the UDCCN handles uplink-downlink calibration for initialization, while the DCEN uses the proposed spatial-frequency sampling embedding module to reduce computational complexity and generative Transformers for temporal dependency learning and autoregressive channel extrapolation.
@@ IV-A Channel Reciprocity Aided Uplink-Downlink Channel Calibration
Although the physical uplink-downlink channels within the channel coherence time are generally reciprocal, uplink-downlink channel calibration is practically necessary to compensate for the hardware asymmetry in transceivers[20]. Specifically, downlink and uplink channels are affected by the transmit and receive hardware (e.g., amplifiers, filters, mixers), respectively. The hardware impairments, such as amplifier nonlinearities, in-phase (I) / quadrature (Q) imbalances, or phase noise, can make the practical downlink and uplink channel responses non-reciprocal. These hardware impairments affect the ideal reciprocity and thus require uplink-downlink channel calibration to adjust for these mismatches. The uplink-downlink channel extrapolation can be formulated by
@@ IV-B Temporal Dependency Learning and Downlink Channel Extrapolation
It is worth mentioning that the downlink channel extrapolation problem is conducted along the time domain and aims to predict channels at future slots, which can be regarded as a time series prediction problem. However, different from general time series prediction problems, historical observations in the investigated downlink channel extrapolation problem (i.e., downlink channels at all downlink slots) can not be obtained unless huge time-frequency resources were spent to conduct channel estimation for all downlink slots. As a result, the performance of widely adopted non-autoregressive prediction schemes in related works[18] would be significantly degraded due to the lack of available historical channel information. Note that like many real-world phenomena, wireless channels exhibit temporal dependencies over time. Then, based on the chain rule in the probability theory, the joint probability of the concatenated downlink channels can be expressed as
@@
where $F_{dce}$ denotes the DCEN, such that (IV-B) can be achieved and the downlink channels from the $2$nd slot to the ($N_{slot}-1$)-th slot can be extrapolated given only $\hat{\mathbf{H}}^{d}\{s,1\}$. It is worth noting that (IV-B) actually refers to the autoregressive generation manner and the DCEN should thus be an autoregressive model. And considering that $\hat{\mathbf{H}}^{d}\{s,t\},t=1,2,\ldots,N_{slot}-1$ are dependent, capturing these temporal dependencies among them would be beneficial to achieve (IV-B). To this end, we propose the DCEN constituted by (inverse) spatial-frequency sampling embedding layers to reduce computational complexity and generative Transformers for temporal dependency learning and autoregressive downlink channel extrapolation. The architecture of the DCEN is also shown in Fig. 2.